40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Mon Jul 26 01:30:42 2021, Koji, Summary, Cryo vacuum chamber, Prep for the 2nd cooling of the suspension 
    Reply  Mon Jul 26 22:49:00 2021, Koji, Summary, Cryo vacuum chamber, Aquadag painting 10x
       Reply  Wed Jul 28 00:59:47 2021, Koji, Summary, Cryo vacuum chamber, The test mass successfully suspended P_20210727_154143.jpegP_20210727_190356.jpegP_20210727_190426.jpegP_20210727_190543.jpeg
          Reply  Thu Jul 29 00:42:38 2021, Koji, Summary, Cryo vacuum chamber, The test mass successfully suspended 
             Reply  Thu Jul 29 13:11:17 2021, Koji, Summary, Cryo vacuum chamber, Cooling progress: Update temp_log_cool_down_20210728_1830.pdfcooling_model1.pdfcooling_model2.pdfOSEM_cooling.pdf
                Reply  Fri Jul 30 12:22:56 2021, Koji, Summary, Cryo vacuum chamber, Cooling curve comparisons comparison_cold_head.pdfcomparison_inner_shield.pdfcomparison_test_mass.pdf
                   Reply  Sun Aug 1 22:22:00 2021, Koji, Summary, Cryo vacuum chamber, Cooling update temp_log_cool_down_20210728_1830.pdfcooling_meas.pdfOSEM_cooling.pdfcooldown_210728.zip
                      Reply  Mon Aug 2 13:25:32 2021, Stephen, DailyProgress, Cryo vacuum chamber, Warmup started 02 August 
                         Reply  Tue Aug 3 23:56:00 2021, Koji, DailyProgress, Cryo vacuum chamber, Warmup started 02 August P_20210804_000247.jpgP_20210803_235421.jpg
Message ID: 2629     Entry time: Sun Aug 1 22:22:00 2021     In reply to: 2625     Reply to this: 2630
Author: Koji 
Type: Summary 
Category: Cryo vacuum chamber 
Subject: Cooling update 

The test mass temperature indicates 121K@100hr but there seemed a few sensor glitches for the test mass (𝛥=-4.2K) and the inner shield (𝛥=-0.43K).
So the actual test mass temperature could be 125K.

The temp was read to be 119K@114hr (Attachment 1)

There was very little cooling capability left for the test mass (Attachment 2)

The OSEM reading is now stable @12.3V (Attachment 3)

The raw temp data and the minimal plotting code are attached (Attachment 4)

Attachment 1: temp_log_cool_down_20210728_1830.pdf  20 kB  Uploaded Tue Aug 3 19:21:13 2021  | Hide | Hide all
temp_log_cool_down_20210728_1830.pdf
Attachment 2: cooling_meas.pdf  151 kB  Uploaded Tue Aug 3 19:21:26 2021  | Hide | Hide all
cooling_meas.pdf
Attachment 3: OSEM_cooling.pdf  26 kB  Uploaded Tue Aug 3 19:22:19 2021  | Hide | Hide all
OSEM_cooling.pdf
Attachment 4: cooldown_210728.zip  193 kB  Uploaded Tue Aug 3 19:22:38 2021
ELOG V3.1.3-